On January 1, 2021, White Water issues $460,000 of 5% bonds, due in 15 years, with interest payable annually on December 31 each year.

Assuming the market interest rate on the issue date is 6%, the bonds will issue at $415,326.

Required:

1. Complete the first three rows of an amortization schedule. (Round your final answers to the nearest whole dollar.)

Date Cash Paid Interest Expense Increase in Carrying Value Carrying Value
01/01/2021
12/31/2021
12/31/2022
Answer #1
Date Cash Paid Interest expense Increase in Carrying Value Carrying value
01/01/2021 415326
12/31/2021 23000 24920 1920 417246
12/31/2022 23000 25035 2035 419281
Workings:
Cash paid 23000 =460000*5%
Interest expense:
12/31/2021 24920 =415326*6%
12/31/2022 25035 =417246*6%